OpenStudy (anonymous):

simplify 15 exponent 18 over 15 exponent 3

OpenStudy (anonymous):

post the multiple choice pls

OpenStudy (anonymous):

A)15 exponent 6 B)15 exponent 15 C)1 exponent 6 D)1 exponent 15

OpenStudy (jtvatsim):

The question is \[\frac{15^{18}}{15^{3}}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its asking me to simplify that

OpenStudy (jtvatsim):

OK, the important thing to remember about exponents is that they usually represent multiplying the same thing over and over. So for example, \[15^3 = 15 \times 15 \times 15\]. where the exponent 3 tells me to multiply 15 three times. Does that make sense as an example?

OpenStudy (anonymous):

yes, thank you but based on the multiply choice what is the answer

OpenStudy (jtvatsim):

Of course, that's the next step. :) I'll help you figure it out by asking you a couple questions. It's not too hard. :) Let's think about the top part of the fraction. How many times do I multiply 15 when I have an exponent of 18?

OpenStudy (anonymous):

18 times

OpenStudy (jtvatsim):

Excellent! So we know that our fraction looks like this...

OpenStudy (anonymous):

:T

OpenStudy (jtvatsim):

\[\frac{15 \times 15 \times \cdots \times 15 \ (18 \text{\times}))}{15 \times 15 \times 15}\]

OpenStudy (jtvatsim):

There are eighteen 15s on top, and three 15s on the bottom. There's going to be cancellations!

OpenStudy (anonymous):

i know

OpenStudy (jtvatsim):

Good, we get to cancel three 15s, which leaves us with fifteen 15s on top.

OpenStudy (marihelenh):

Dividing the two is also equal to\[15^{18-3}\] It will get you the same answer.

OpenStudy (jtvatsim):

The final answer is that we have fifteen 15s multiplied together. This is the same as \[15^{15}\] we are done!

OpenStudy (anonymous):

Thanks

OpenStudy (jtvatsim):

Your welcome, good job!